# Cross-border Exchange of CBDCs using Layer-2 Blockchain

### 22 Dec 2023 v1

@@ Abstract
This paper proposes a novel multi-layer blockchain architecture for the cross-border trading of CBDCs. The permissioned layer-2, by relying on the public consensus of the underlying network, assures the security and integrity of the transactions and ensures interoperability with domestic CBDCs implementations. Multiple Layer-3s operate various Automated Market Makers (AMMs) and compete with each other for the lowest costs. To provide insights into the practical implications of the system, simulations of trading costs are conducted based on historical FX rates, with Project Mariana as a benchmark. The study shows that, even with liquidity fragmentation, a multi-layer and multi-AMM setup is more cost-efficient than a single AMM.
@@
Decentralized Finance Automated Market Maker CBDC Layer-2 Blockchain
@@ 1 Introduction
The Bitcoin network, launched in 2008, was first described in a whitepaper by the anonymous individual or group, known as Satoshi Nakamoto [1]. The fundamental promise of this first cryptocurrency was to establish a peer-to-peer payment system that operates independently of traditional financial intermediaries. In 2014, Vitalik Buterin introduced Ethereum [2], the next generation blockchain with a virtual machine capable of executing computer programs known as smart contracts. With Ethereum and the Ethereum Virtual Machine (EVM), Decentralized Finance (DeFi) was born. By utilizing smart contracts, DeFi offers sophisticated financial services, such as trading, lending, borrowing, derivatives and asset management [3, 4, 5, 6].
It quickly became apparent that this new blockchain-based financial system required non-volatile tokens, giving the advent to stablecoins—tokens with values pegged to fiat currencies. The success of stablecoins prompted central banks to explore similar solutions, leading to the development of Central Bank Digital Currencies (CBDCs) [7]. While CBDCs do not necessarily have to be implemented on a blockchain, the decentralized approach promises enhanced scalability and security [8, 9, 10, 11]. Various central banks across the globe are already engaged in CBDC initiatives[8, 9, 12]. The collective effort explores ways to integrate these CBDCs into a unified marketplace, potentially transcending borders, as seen in Project Mariana by the Bank for International Settlements (BIS) [8].
This paper introduces a novel architectural framework for cross-border CBDC trading that is based on Layer-2 (L2) blockchain scaling. L2s proved to be an efficient approach for addressing the limitations of Ethereum, especially scalability and privacy. [13, 14, 15]. The proposed system is the rollup on the public L1 blockchain that seamlessly integrates additional Layer-3 (L3) blockchains, each operating decentralized exchanges (DEXes), or other DeFi protocols, utilizing CBDCs. The system automatically selects L3 and DEXes with optimal costs. This system architecture is compared to the approach of Project Mariana in a series of simulations based on the historical Foreign Exchange (FX) rates.
@@ 2 Background
This section introduces the concept of Layer-2 blockchain, presents the functionalities of Project Mariana, and summarizes this work’s contributions to the research on cross-border CBDCs exchange.
@@
2.1 Layer-2 Blockchains
There are two approaches to tackle the challenges of blockchain scalability and related high gas prices: Layer-1 and Layer-2 blockchains. Layer-1 scaling involves creating entirely new blockchains, e.g., with unique consensus mechanisms or block sizes. These chains operate with their own validators and infrastructure, requiring decentralization to win users’ trust. Layer-2 scaling takes a different approach and uses complex computations off-chain (outside of Ethereum), aiming to reduce on-chain data congestion. These off-chain activities include rollups, sidechains, plasma, and state channels. Layer-2 blockchains rely on the security of underlying Layer-1 chains, such as Ethereum, for final settlement.
Rollups represent a form of Layer-2 scaling that does not custody any data by themselves. Rollups offload complex calculations from the Ethereum mainnet and store the results (along with other transactions) in Ethereum after compressing them. Fig. 2 illustrates rollup architecture with key components—sequencers and verifiers. Sequencers roll up transactions to the Layer-1 chain. By bundling transactions, rollups manage to save on gas fees. Verifiers are smart contracts that operate on Ethereum and verify the transactions stored by the sequencer. They ensure the correctness of the transactions.
2.2 Project Mariana
The specific focus of this work is on facilitating cross-border exchanges of CBDCs. One notable proposal for such a solution is Project Mariana (L1-Mariana) [8], initiated by BIS, the Swiss National Bank, the Bank of France, and the Monetary Authority of Singapore.
Within L1-Mariana, a DEX with an Automated Market Maker (AMM) is employed to exchange wholesale CBDCs versions of Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD). The selected AMM is the Stableswap Invariant, first introduced by Curve v2—the second largest DEX in terms of trading volumes in permissionless DeFi on Ethereum [16]. The setup involves bridges that facilitate the transition of wholesale CBDCs from domestic central bank blockchains to the chain where the Cryptoswap Invariant pool operates, as depicted in Fig. 2.
@@ 3 Related Work and Contribution
There is an abundance of literature discussing the implementation of CBDCs in various countries [9, 10, 11, 17, 12], but little research on the exchange between CBDCs in AMMs, except for Project Mariana [8]. Lipton and Sepp[18] study the implicit costs (slippage cost) for G-10 CBDCs exchange at AMM with Concentrated Liquidity (Uniswap v3).
Our work is the first research that takes the holistic approach to analyze the total costs of CBDCs exchange at various AMMs, including Stableswap Invariant (Curve v2), applied in Project Mariana. This study found that a multi-AMM set-up for CBDCs exchange on L2 is more cost-efficient than a single AMM of Project Mariana, despite the fragmentation of liquidity among various AMMs. Such a multi-AMM system on private L2 is further presented, and its behavior is evaluated via simulation based on the historical FX rates, Mariana Project set up as a benchmark.
@@
[…]
@@ 4 System Architecture
Figure 3: General Architecture of Cross-Border CBDC Exchange using L3 and L2s on Public Layer-1 Network
@@
5 Metrics and AMMs
The major metric to consider for evaluation of the cross-border CBDCs exchange is the cost of the CBDCs swap. In general, the total cost of DeFi service consists of three components: [23, 24]
@@ 5 Metrics and AMMs
Gas Fees are paid to the blockchain operators for executing the transaction, and DeFi Explicit Fee to the DeFi protocol for the provided service. DeFi Implicit Fee are indirect fees specific to the design of the DeFi protocol. For tokens swaps at AMM-DEX, swap fees is the explicit fee charged by the DEX, whereas slippage (price impact) is the implicit fee: [23, 24]
$TotalSwapCost=GasFee+SwapFee+SlippageCost$ (2)
SwapFee is typically the percentage of the transaction volume, and Liquidity Providers (LPs) supply capital to the AMM’s liquidity pools in exchange for participation in swap fees. Slippage Cost is the change in exchange price caused by the transaction and is specific to the AMM type, the liquidity pool size and transaction volume. The next section presents the major AMM types and corresponding slippage calculation methods.
@@ 5.1 Automated Market Makers
Cryptoswap Invariant - Curve v2  As in equation (4), but with K defined as
@@ 5.2 Slippage at AMM-DEXs
$Slippage(\Delta x,\Delta y,S_{t})=\frac{\Delta y/\Delta x}{S_{t}}-1$ (7)
@@
$Slippage(t)=\frac{\Delta y}{N_{0}*\sqrt{S_{0}*S_{t}}}(1-\frac{1}{\sqrt{\alpha}})$ (8)
@@ 6 Simulation
To ensure comparability with L1-Mariana , we consider three CBDCs: Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD) and back-test the total costs of CBDCs swaps in L1-Mariana and L2-L3-Exchange systems. We consider the buy transaction of EUR for CHF for volumes ranging from 1 EUR to 1mn EUR. The slippage costs for CLMMs are calculated based on Eq. 8, and the Newton-Raphson method is applied for Cryptoswap Invariants, both for $n=2$ and $n=3$. The implementation of all methods is available in a public GitHub repository [19] and was based on the smart contract code of Curve v2.
@@ 6.1 Data and Parameters
Gas costs.  Gas fees for swaps on Ethereum, referred to in the Project Mariana, oscillated between 5 EUR and 50 EUR over the last three years [29]. The gas costs on L2s are, on average, 50 times lower compared to the underlying L1 [15, 14]. Based on these facts, we consider 15 EUR gas fees for a swap on L1 and 15EUR/50 on the L2-L3 network.
@@ 6.2 Results
Figure 4: Difference between the total swap costs at L1-Mariana and L2-L3-Exchange. The negative difference indicate L2-L3-Exchange out-performance for small and large transaction volumes.
Figure 5: Breakdown of total swap costs for various transaction volumes
Figure 6: Slippage costs for L1 Mariana (100mn TVL) and L3 AMM operators (each 33mn TVL)
The simulation outcomes are depicted in Figs. 4, 5 and 6. Fig. 8 in the appendix presents the historical total swap costs for transaction volumes of 10k, 100k, and 1mn EUR for CHF for each day over the last 3 years with fixed gas fees. Fig 4 extends this analysis for various gas fees and presents the difference between the costs of L2L3 Exchange and L1 Mariana. Notably, L2-L3-Exchange outperforms for volumes of 10k and 1mn. This out performance varies, influenced not only by transaction volume but also by current gas fees, as shown in fig. 4. For gas fees surpassing 800 EUR, L2-L3-Exchange outperforms for any transaction volume. For lower gas fees, such as 15 EUR, the out performance range is observed for volumes representing less than 0.01% or more than 0.7% of the provided liquidity. The reason for L2-L3-Exchange outperforming for smaller transactions lies in the minimization of gas fees through the application of L2 and L2s. Fig. 5 breaks down total swap costs, indicating that gas fees dominate the costs for L1 Mariana in smaller transactions, compared to the swap fee for L2-L3-Exchange. For larger transactions, slippage becomes the predominant component of total swap costs.
The out performance of the L2-L3-Exchange over L1-Mariana and its ability to handle slippage effectively stem from the use of L3 with CLMM, as illustrated in Figs. 6. The analysis of L3 selection indicates the use of L3 with Cryptoswap Invariant 2-token pool for smaller transactions, while the 3-token pool of Cryptoswap Invariant is unused for any transaction volumes. The simulation showed that the application of more AMMs, despite the liquidity fragmentation, outperforms the application of just one AMM: 100mn to L1-Mariana and 33.33mn to each L3. Fig. 9 simulates the swap costs when various total liquidity is provided.
@@ 7 Discussion
Furthermore, the profitability of liquidity provisions and corresponding swap fee, should be analyzed. The prototype implementation of such systems could be based on the ready rollup frameworks [30, 31, 32].

### 30 Jan 2024 v2

@@ Abstract
This paper proposes a novel multi-layer blockchain architecture for the cross-border trading of CBDCs. The permissioned layer-2, by relying on the public consensus of the underlying network, ensures the security and integrity of the transactions and interoperability with domestic CBDC implementations. Multiple Layer-3s operate various Automated Market Makers (AMMs) and compete with each other for the lowest costs. Simulations of trading costs are conducted based on historical FX rates to provide insights into the practical implications of the system, with Project Mariana as a benchmark. The study shows that a multi-layer and multi-AMM setup is more cost-efficient even with liquidity fragmentation than a single AMM.
@@
Decentralized Finance Automated Market Maker
Central Bank Digital Currency Blockchain
@@ 1 Introduction
The Bitcoin network, described in a whitepaper by the anonymous individual or group known as Satoshi Nakamoto, launched in 2008 [22]. The fundamental promise of this first cryptocurrency was to establish a peer-to-peer payment system that operates independently of traditional financial intermediaries. In 2014, Vitalik Buterin introduced Ethereum [10], the next-generation blockchain with a virtual machine capable of executing computer programs known as smart contracts. With Ethereum and the Ethereum Virtual Machine (EVM), Decentralized Finance (DeFi) was born. By utilizing smart contracts, DeFi offers sophisticated financial services, such as trading, lending, borrowing, derivatives and asset management [29], [25], [9], [18].
It quickly became apparent that this new blockchain-based financial system required non-volatile tokens, giving the advent to stablecoins—tokens with values pegged to fiat currencies. The success of stablecoins prompted central banks to explore similar solutions, leading to the development of Central Bank Digital Currencies (CBDCs) [28]. While CBDCs do not necessarily have to be implemented on a blockchain, the decentralized approach promises enhanced scalability and security [4], [23], [13], [12]. Various central banks across the globe are already engaged in CBDC initiatives[4], [23], [5]. The collective effort explores ways to integrate these CBDCs into a unified marketplace, potentially transcending borders, as seen in Project Mariana by the Bank for International Settlements (BIS) [4].
This paper introduces a novel architectural framework for cross-border CBDC trading that is based on Layer-2 (L2) blockchain scaling. L2s proved to be an efficient approach for addressing the limitations of Ethereum, especially scalability and privacy [26], [32], [16]. The proposed system is the rollup (non-custodial L2), on the public L1 blockchain that seamlessly integrates additional Layer-3 (L3) blockchains, each operating decentralized exchanges (DEXes), or other DeFi protocols, utilizing CBDCs. The system automatically selects L3 and DEXes with optimal costs. This system architecture is compared to the approach of Project Mariana in a series of simulations based on the historical Foreign Exchange (FX) rates.
@@ 2 Background
This section introduces the concept of L2 blockchain scaling, presents the functionalities of Project Mariana, and summarizes this work’s contributions to the research on cross-border CBDCs exchange.
2.1 Layer-2 Blockchains
There are two approaches to tackle the challenges of blockchain scalability and related high gas prices: Layer-1 (L1) and Layer-2 (L2) blockchains. L1 scaling involves creating entirely new blockchains, e.g., with unique consensus mechanisms or block sizes. These chains operate with their own validators and infrastructure, requiring decentralization to win users’ trust. L2 scaling takes a different approach and uses complex computations off-chain (outside of Ethereum), aiming to reduce on-chain data congestion. These off-chain activities include rollups, sidechains, plasma, and state channels. Layer-2 blockchains rely on the security of underlying Layer-1 chains, such as Ethereum, for final settlement.
@@ 2.1 Layer-2 Blockchains
Rollups represent a form of L2 scaling that does not custody any data by themselves. Rollups offload complex calculations from the Ethereum mainnet and store the results (along with other transactions) in Ethereum after compressing them. Fig. 1 illustrates rollup architecture with key components—sequencers and verifiers. Sequencers roll up transactions to the Layer-1 chain. By bundling transactions, rollups manage to save on gas fees. Verifiers are smart contracts that operate on Ethereum and verify the transactions stored by the sequencer. They ensure the correctness of the transactions.
2.2 Project Mariana
The specific focus of this work is on facilitating cross-border exchanges of CBDCs. One notable proposal for such a solution is Project Mariana on L1 blockchain (L1-Mariana) [4], initiated by BIS, the Swiss National Bank, the Bank of France, and the Monetary Authority of Singapore.
@@ 2.2 Project Mariana
Within L1-Mariana, a DEX based on AMM is employed to exchange wholesale CBDCs versions of Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD). The selected AMM is the Stableswap Invariant, first introduced by Curve v2—the second largest DEX in terms of trading volumes in permissionless DeFi on Ethereum [1]. The setup involves bridges that facilitate the transition of wholesale CBDCs from domestic central bank blockchains to the chain where the Cryptoswap Invariant pool operates, as depicted in Fig. 2.
@@ 3 Related Work and Contribution
There is an abundance of literature discussing the implementation of CBDCs in various countries [23], [13], [12], [21], [5], but little research on the exchange between CBDCs in AMMs, except for Project Mariana [4]. Lipton and Sepp[20] study the implicit costs (price impact) for G-10 CBDCs exchange at AMM with Concentrated Liquidity (Uniswap v3).
This work is the first research that takes the holistic approach to analyze the total costs of CBDCs exchange at various AMMs, including Stableswap Invariant (Curve v2), applied in Project Mariana. This study found that a multi-AMM set-up for CBDCs exchange on L2 is more cost-efficient than a single AMM of Project Mariana, despite the fragmentation of liquidity among various AMMs. Such a multi-AMM system on private L2 is further presented, and its behavior is evaluated via simulation based on the historical FX rates, Mariana Project set up as a benchmark.
@@
  • Demonstrating that multi-AMMs set-up for CBDCs exchange is more cost-efficient than a single AMM.
  • Designing such multi-AMMs set-up for CBDCs exchange using L2 and L3 blockchains and demonstrating its superior performance for small, medium, and largest transaction volumes and any transaction in case of spikes of gas fees.
  • Conducting first detailed analysis of the total costs of CBDCs swaps on AMM-DEX—gas fees, swap fees (explicit fees), and price impact (explicit fees), among various AMMs and L1 and L2-based DEXes.
  • Providing quantitative metrics to evaluate the efficiency of cross-border CBDCs on DEXes for major AMMs. Their implementation is available in the public GitHub repository [17]
  • Building a framework for simulation based on historical FX rate that can be replicated in other cross-border CBDCs projects or in permission DeFi in the Ethereum ecosystem. For all metrics, the mathematical background and numerical methods with the implementation are provided, such as the Newthon-Raphon method for price impact in Cryptoswap Invariant AMM applied in Project Mariana.
@@ 4 System Architecture
Figure 3: General Architecture of Cross-Border CBDCs Exchange using L3 and L2 Blockchains on Public L1 Blockchain
@@
5 Model
The major metric to consider for evaluation of the cross-border CBDCs exchange is the cost of the CBDCs swap. In general, the total cost of DeFi service consists of three components: [31], [19]
@@ 5 Model
Gas Fees are paid to the blockchain operators for executing the transaction. In the case of transactions on roll-ups, the gas fee is comprised of both the L1 and L2 gas fee components.
DeFi Explicit Fee is charged by the DeFi protocol for the provided service. In the case of token swap at AMM-DEX, the liquidity pool fee (LP Fee) is the explicit fee charged by the DEX, and part of it (or whole) is distributed to the liquidity providers (LPs) to the pool.
DeFi Implicit Fee are indirect fees specific to the design of the DeFi protocol. Price impact and slippage are implicit fees accrued by a trader when swapping tokens at AMM [31], [19] and refer to the difference between quoted and executed price by the AMM. Price impact is caused by the transaction volume and slippage (positive or negative) results from other swaps occurring within the same block as the transaction swap.
Further the following model for the total cost of swap is considered:
$TotalSwapCost=GasFee+LPFee+PriceImpact$ (2)
LP Fee is typically the percentage of the transaction volume, and liquidity providers supply capital to the AMM’s liquidity pools in exchange for participation in LP fees. Price impact is the change in exchange price caused by the transaction volume and is specific to the AMM type and the liquidity pool size. The next section presents the major AMM types and corresponding price impact calculation methods.
@@ 5.1 Automated Market Makers
Cryptoswap Invariant - Curve v2
As in equation (4), but with K defined as
@@ 5.2 Price Impact
$PriceImpact(\Delta x,\Delta y,S_{t})=\frac{\Delta y/\Delta x}{S_{t}}-1$ (7)
@@
$PriceImpact(t)=\frac{\Delta y}{N_{0}*\sqrt{S_{0}*S_{t}}}(1-\frac{1}{\sqrt{\alpha}})$ (8)
@@ 6 Simulation
To ensure comparability with L1-Mariana , we consider three CBDCs: Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD) and back-test the total costs of CBDCs swaps in L1-Mariana and L2-L3-Exchange systems. We consider the buy transaction of EUR for CHF for volumes ranging from 1 EUR to 1mn EUR. The price impact for CLMMs are calculated based on Eq. 8, and the Newton-Raphson method is applied for Cryptoswap Invariants, both for $n=2$ and $n=3$. The implementation of all methods is available in a public GitHub repository [17] and was based on the smart contract code of Curve v2.
@@ 6.1 Data and Parameters
Gas costs. Gas fees for swaps on Ethereum, referred to in the Project Mariana, oscillated between 5 EUR and 50 EUR over the last three years [6]. The gas costs on L2s are, on average, 50 times lower compared to the underlying L1 [16], [32]. Based on these facts, we consider 15 EUR gas fees for a swap on L1 and 15EUR/50 on the L2-L3 network.
@@ 6.2 Results
Figure 4: Difference between the total swap costs at L1-Mariana and L2-L3-Exchange. The negative difference indicates L2-L3-Exchange out-performance for small and large transaction volumes.
Figure 5: Breakdown of total swap costs into gas fees, LP fees (swap fee) and price impact (slippage) as a function of the swap volume
Figure 6: Price impact at L1 Mariana and L3 AMM operators. The liquidity pool at L1 Mariana AMM is worth 100mn EUR and each pool of L3 AMM is worth 33mn EUR
The simulation outcomes are depicted in Figs. 4, 5 and 6. Fig. 8 in the appendix presents the historical total swap costs for transaction volumes of 10k, 100k, and 1mn EUR for CHF for each day over the last 3 years with fixed gas fees. Fig 4 extends this analysis for various gas fees and presents the difference between the costs of L2L3 Exchange and L1 Mariana. Notably, L2-L3-Exchange outperforms for volumes of 10k and 1mn. This outperformance varies, influenced not only by transaction volume but also by current gas fees, as shown in fig. 4. For gas fees surpassing 800 EUR, L2-L3-Exchange outperforms for any transaction volume. For lower gas fees, such as 15 EUR, the out performance range is observed for volumes representing less than 0.01% or more than 0.7% of the provided liquidity. The reason for L2-L3-Exchange outperforming for smaller transactions lies in the minimization of gas fees through the application of L2 and L2s. Fig. 5 breaks down total swap costs, indicating that gas fees dominate the costs for L1 Mariana in smaller transactions, compared to the swap fee for L2-L3-Exchange. For larger transactions, price impact becomes the predominant component of total swap costs.
The outperformance of the L2-L3-Exchange over L1-Mariana and its ability to handle price impact effectively stem from the use of L3 with CLMM, as illustrated in Figs. 6. The analysis of L3 selection indicates the use of L3 with Cryptoswap Invariant 2-token pool for smaller transactions, while the 3-token pool of Cryptoswap Invariant is unused for any transaction volumes. L3 with CLMM and 2-token pool Cryptoswap Invariant outperforms, in certain situations, the 3-token pool of CryptoSwap Mariana despite liquidity fragmentation, with each L3 having three times less liquidity than L1 Mariana. Fig. fig. 5 explains the role of gas, swap, and price impact components in overall transaction costs. The simulation showed that applying more AMMs, despite the liquidity fragmentation, outperforms the application of just one AMM: 100mn to L1-Mariana and 33.33mn to each L3. Fig. 9 simulates the swap costs when various total liquidity is provided.
Table 1: Breakdown of the average fees for the small (10k EUR), medium (100k EUR) and large (1mn EUR) transactions. The buy EUR-CHF transactions are considered at L1 Mariana and L2-L3 Exchange. All fees are denominated in EUR. The simulation is based on the daily historical FX rate for the last three years.
Total Fee Gas Fee Swap Fee Price Impact
Size
L1 Mariana:
Small 16.0 (16bps) 15.0 (15bps) 1.0 (1bps) 0.01 (0bps)
Medium 26.1 (3bps) 15.0 (2bps) 10.0 (1bps) 1.08 (0bps)
Large 10,386.4 (104bps) 15.0 (0bps) 100.0 (1bps) 10,271.42 (103bps)
L2 Exchange:
Small 1.3 (1bps) 0.3 (0bps) 1.0 (1bps) 0.02 (0bps)
Medium 25.6 (3bps) 0.3 (0bps) 10.0 (1bps) 15.33 (2bps)
Large 5,445.3 (54bps) 0.3 (0bps) 100.0 (1bps) 5,345.04 (52bps)
@@ 7 Discussion
Furthermore, the profitability of liquidity provisions and corresponding swap fee, should be analyzed. The prototype implementation of such systems could be based on the ready rollup frameworks [3], [2], [7].